# NuSegDG: Integration of Heterogeneous Space and Gaussian Kernel for Domain-Generalized Nuclei Segmentation

### v2

@@ 4.4 Comparison on Nuclei Instance Segmentation
4.5 Ablation Study
To investigate the effectiveness of the individual components within the NuSegDG framework, we conduct an ablation study on domain-generalized nuclei instance segmentation, as summarized in Table 5. This study sequentially enables or disables the HS-Adapter $M_{1}$, GKP-Encoder $M_{2}$, and TSM-Decoder $M_{3}$ to evaluate their impact on the performance of the average AJI, DQ, SQ, and PQ metrics. Firstly, we consider the standard fine-tuned SAM ($1^{st}$ row) as the ablation baseline. By respectively embedding the HS-Adapter ($2^{nd}$ row), GKP-Encoder ($3^{rd}$ row) and TSM-Decoder ($4^{th}$ row), the performance is raised with the average AJI of 5.17%, 2.79%, 2.07%, and the average PQ of 6.61%, 3.17%, 1.58%.
Moreover, when we combine HS-Adapter with GKP-Encoder ($5^{th}$ row), the performance of the model is further improved, with the average AJI of 56.29 and PQ of 50.38% on the four domains. This result proves that these two modules can promote the domain generalization capability in nuclei instance segmentation. By comparing $6^{th}$ and $7^{th}$ rows with $2^{nd}$ and $3^{rd}$ rows, the TSM-Decoder demonstrates significant performance gains while eliminating the demand for manual morphological refinement. Finally, our NuSegDG framework ($8^{th}$ row) integrates all three modules and achieves the best performance on all metrics, with an average AJI of 58.84%, an average DQ of 69.62%, an average SQ of 75.50%, and an average PQ of 53.21%. This full configuration significantly outperforms the others, emphasizing the synergistic benefits of incorporating all modules. This result highlights the importance of each component in enhancing the generalization capability of NuSegDG across different nuclei image domains.
4.6 Analysis of Hyper-Parameters
In this section, we perform a comprehensive hyper-parameters analysis of our NuSegDG model. As reported in Section 3.2 and 3.3, NuSegDG contains two hyper-parameters, including the Gaussian kernel size $r$ in GKP-Encoder and the number of heterogeneous space $N$ in HS-Adapter. For the kernel size, we perform a grid search under the fully-supervised learning to select an optimal configuration. Fig. 4a shows the average Dice and mIoU of NuSegDG on the four nuclei domains with different kernel sizes. It is indicated that the NuSegDG with $r=10$ demonstrates the best performance due to the sufficient semantic information prompts. However, excessive kernel size may generate false positive errors, which cannot offer additional benefits. For the number of heterogeneous space, we provide the result of grid search in Fig. 4b. We observe that the NuSegDG with $N=2$ obtains the best performance. Setting more heterogeneous space significantly increases the computational complexity of NuSegDG, which is not suitable for limited training samples in nuclei domains. These experimental results prove the importance of tuning these hyper-parameters to improve the efficiency of our NuSegDG framework in learning domain-specific knowledge.
@@ 5 Conclusion
In this paper, we have proposed NuSegDG for domain-generalized nuclei image segmentation. Specifically, the HS-Adapter has been introduced to adapt the feature representation of SAM from natural to different nuclei images by heterogeneous space integration. Then, the GKP-Encoder has been devised to produce high-quality density maps, driven by the single-point prompt, with sufficient semantic information for guiding segmentation predictions. Finally, the TSM-Decoder has achieved the automatic conversion between the semantic masks and instance maps without demand for labor-intensive morphological post-processing methods. Extensive experimental results have demonstrated that NuSegDG has outperformed the existing nuclei-specific and SAM-based segmentation methods in domain-generalized nuclei image segmentation and displayed superior adaptability across different nuclei domains.

### v3

@@ 5 Conclusion
In this paper, we have proposed NuSegDG for domain-generalized nuclei image segmentation. Specifically, the HS-Adapter has been introduced to adapt the feature representation of SAM from natural to different nuclei images by heterogeneous space integration. Then, the GKP-Encoder has been devised to produce high-quality density maps, driven by the single-point prompt, with sufficient semantic information for guiding segmentation predictions. Finally, the TSM-Decoder has achieved the automatic conversion between the semantic masks and instance maps without demand for labor-intensive morphological post-processing methods. Extensive experimental results have demonstrated that NuSegDG has outperformed the existing nuclei-specific and SAM-based segmentation methods in domain-generalized nuclei image segmentation and displayed superior adaptability across different nuclei domains. The proposed NuSegDG presents a potential nuclei annotation tool for improving the efficiency of data labeling, and its accurate delineation of nuclei can aid in tumor detection, grading, and diagnostic assessments.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
This work is partially supported by the Yongjiang Technology Innovation Project (2022A-097-G), Zhejiang Department of Transportation General Research and Development Project (2024039), and National Natural Science Foundation of China grant (UNNC ID: B0166).